OpenStudy (anonymous):

Find the value of x that makes the statement true. 6 + 5.6 – x = 6 A. –6 B. 6 C. 5.6 D. –5.6

OpenStudy (texaschic101):

6 + 5.6 - x = 6 11.6 - x = 6 11.6 - 6 = x 5.6 = x